Can someone check my work to see if I got it right? Thanks
Graph the equation by plotting points. y-3x=0
y-3(0)=0
y=0
(0,0)
0-3x=0
-3x=0
(0,0)


Answer by Edwin McCravy(20081)

The intercept method fails sometimes and you have to get a point that is
not an intercept.  That is the case with your equation. It takes two
points to determine a line, not just one.

There is only one point that is on both the x-axis and the y-axis, and that
is the origin (0,0).  Sometimes the y-intercept and the x-intercept of a line
are both the same point (0,0).  This is what happens with your equation.

Since the origin (0,0) is both the x-intercept as well as the y-intercept, 
that's not enough to draw the graph of a line.  You must find another point
on that line which is neither the x-intercept nor the y-intercept. To do that,
choose any number other than 0 for x and substitute that value for x, and
solve for the corresponding value of y.

Suppose we choose x=1. (We could choose any other non-zero number as well). 
Then we substitute it into the equation and solve for y:

  y - 3x = 0
y - 3(1) = 0
   y - 3 = 0
Add 3 to both sides
       y = 3

So a second point on the line, which is neither an x-intercept nor a y-
intercept is (1,3), so now we have two points, (0,0) and (1,3), and we
can draw the line through two points.

So plot those:

and draw a line thru them:
  
Edwin